OpenStudy (anonymous):

How many moles of calcium chloride would be produced by the reaction below when 53.2 milliliters of water are produced, if the density of water is 0.987 g/mL? Show all steps of your calculations as well as the final answer. HCl + Ca(OH) 2 → H2O + CaCl2

OpenStudy (anonymous):

@shadowferrara Post it in Chemistry section Don't forget to balance, though :) Ca(OH)2+ 2HCl → CaCl2 + 2H20

OpenStudy (anonymous):

sorry about that, i didnt realize...

OpenStudy (anonymous):

Mass of water: m = d* v = .987 * 53.2 = ... g -> Mole of water: = ...g/ 18 = .... mol => Mole of CaCl2 = ... mol / 2

OpenStudy (anonymous):

thank you so much, eventhough it was in the wrong section....

OpenStudy (anonymous):

Can you fill in the blanks to get the result?

OpenStudy (anonymous):

Yes, i believe it is Ca(OH)2+ 2HCl → CaCl2 + 2H20 Mass of water = .987 * 53.2 = 52.5084 g Mole of water = 52.5084g/ 18= 2.917 mol Mole of CaCl2= 2.917 mol / 2 = 1.458

OpenStudy (anonymous):

Excellent =) Most of person in Math section know Chem, but I'm not expert!

OpenStudy (anonymous):

Thats awesome, because it seems there are not that many people in the chem section. If you have time is this right? How many grams of aluminum chloride must decompose to produce 78.3 milliliters of aluminum metal, if the density of aluminum is 2.70 g/mL? Show all steps of your calculation as well as the final answer. AlCl3 → Al + Cl2 2AlCl3 -> 2Al + 3Cl2 2:2 Mass of aluminum = 78.3*2.70=211.41g Moles of aluminum = 211.4g/26.98 = 7.835 grams of alcl= 7.835* (26.98+35.45)= 246.85 g

OpenStudy (anonymous):

Almost, except Molar mass of AlCl3: 26.98 + 3 * 35.45 = ... Mass of AlCl3 : 7.835 * .... =

OpenStudy (anonymous):

I assume calculate is your part, and you also well understand Sig fig. rule :)

OpenStudy (anonymous):

Molar mass of AlCl3: 26.98 + 3 * 35.45 = 133.33 Mass of AlCl3 = 7.835 * 133.33 = 1044.64g yes and yes thank yous for the help!